one thousand dollars in a savings account pays 7% interest per year. The interest earned after the first year is added to the account. How much interest is earned on the new principle the following year?
Thanks you very much

Answer by Nate(3500)

I = PRT
I = 1000*0.07*1 = 70
now ...
I = PRT = 1070*0.07*1 = 74.9
or
P(a) = a(1 + r)^t = 1000(1.07)^2 = 1144.9
Throughout the time, the total interest is $144.9